# Error Estimation for Adaptive Mesh Refinement in Droplet Simulations

### v1

@@
Abstract: We present a one-dimensional shear force driven droplet formation model with a flux-based error estimation. The presented model is derived using asymptotic expansion and a front-tracking method to simulate the droplet interface. The model is then discretized using the Galerkin finite element method in the mixed form. However, the jumps in the solution gradients are discontinuous and can grow faster due to the highly convective pinch-off process. This leads to an erroneous droplet interface and incorrect curvature. Therefore, the mesh must be sufficiently refined to capture the interface accurately. The mixed form of the governing equation naturally provides smooth interface gradients that can be used to compute the error estimate. The computed error estimate is then used to drive the adaptive mesh refinement algorithm.
@@ 1 Introduction
The formation of droplets and bubbles is a mature field in physics. These surface tension driven phenomena have been deeply explored using interface-tracking and interface-capturing approaches. Historically, the accurate description of free surface flows emerged after Laplace [1] and Young [2] explained the role of mean curvature and the surface tension force as a source of instability in 1805. Eggers and Villermaux presented a comprehensive review on understanding the liquid break-up process [3]. Starting with the initial fluid column, the drop becomes heavier by continuously adding fluid. Under an external force like gravity, the instabilities start to grow on the interface. Meanwhile, the surface tension minimizes the surface energy by decreasing the radius of the fluid column. A spherical droplet starts to form at the end of the fluid column hanging from a very thin neck. Eventually, the radius becomes zero and a droplet separates from the initial fluid column. The point of zero radius is called the “pinch-off" point. After the primary pinch-off, the neck recoils due to the “unbalanced” surface tension. The capillary waves due to this recoil perturb the interface before the tip can collapse back to the top of the fluid cone. This leads to a secondary breakup creating smaller droplets, which are called “satellite drops”.
@@
Inaugural work on the mathematical modeling of droplet formation was done by Eggers and Dupont [4]. Their work involved finite difference formulation of the one-dimensional asymptotic model. Ambravaneswaran et al. [5] explored a finite element version of the same model. This model was only explored with the gravitational force as an external force. However, many industrial applications like atomization and droplet entrainment in annular flow [6, 7], and microfluidics devices [8, 9, 10] involves of droplet formation under a shear force environment. Using the same asymptotic approach for shear-induced droplets, we considered a one-dimensional model that estimates droplet sizes for an environment with co-flowing fluids [11]. For applications with a symmetrical structure of droplet formation, this one-dimensional model with the interface tracking approach is advantageous in terms of computational costs, while still maintaining the accuracy required for comparison with experiments. Nonetheless, this shear-induced droplet model deals with a wide range of external force magnitudes that induces a wide range of interface instability time frame.
@@
The instability leading to the pinch-off point requires a finer mesh in the region of high curvature. As the interface evolves with the progression of numerical simulation, the mesh must be adaptively refined to control discretization errors, especially in the singularity regions. Therefore, a robust error estimation method is essential in designing an adaptive mesh refinement algorithm for our one-dimensional model. The field of a posteriori error estimation has been under continuous and rigorous improvement since the 1980s [12]. The preliminary work on error estimation was based on the finite element residuals and the flux jumps across the element boundaries [13, 14, 15, 16]. Since this work, there are many thorough review articles and books written on the topic of error estimation in finite element analysis from which we wish to refer here [17, 18, 19]. A thorough review on the recent advancements in the field of error estimation is in [20]. In the present work, we show that the mixed finite element formulation of our one-dimensional model naturally provides an effective error estimate for adaptive mesh refinement.
@@
In the following sections, we describe the one-dimensional droplet pinch-off model and explain the need for an error estimator. Then, we explore the flux-based error estimation approach and show that the mixed form naturally provides a smooth flux to compute error bounds. Finally, we illustrate the usefulness of our error estimate using adaptive mesh refinement.
@@ 2 Problem Description
We consider a fluid column slowly flowing with the constant velocity $u^{d}$ out of a nozzle with the radius $h_{in}$. The continuous (or outer) fluid is co-flowing with the dispersed (or inner) fluid column but with a higher velocity $u^{c}$ in a capillary tube with radius $R$. Fig. (1) shows the schematic describing this co-flowing fluids scenario. The superscripts $d$ and $c$ represent dispersed phase and continuous phase fluids, respectively.
@@
We start with the Navier-Stokes equations in the cylindrical coordinates. The pinch-off process happens in finite time due to the surface tension forces trying to minimize the surface energy by contracting the interface in the radial direction. The radial contraction is faster than the axial elongation, which allows us to expand the solution variables asymptotically in $r$. Finally, the normal forces are balanced by the surface tension forces, which are used to simplify the equation into a one-dimensional model.
Similarly, we consider the droplet pinch-off in another co-flowing fluid under the gravitational and shear forces. We consider the co-flowing fluids with a preserved symmetry around the angular direction. This provides the basis for the one-dimensional approach. The continuous fluid affects the droplet morphology by applying shear force on the interface between two fluids. This shear force effect can be approximated with the asymptotically derived force balance on the interface, leading to the derivation of the one-dimensional governing equations [11]. The full derivation can be found in [21].
@@
Here, parameters $\gamma$, $\rho$, $\nu$, $\mu$, and $p$ represent surface tension coefficient, density, kinematic viscosity, dynamic viscosity, and pressure, respectively. The superscript represents dispersed or continuous phase fluid. $\mathcal{K}$ is the curvature defined by
@@
The quantity (C1)h(C-1)h defines the thickness of the shear layer in the continuous phase flow, which defines how much force is experienced by the dispersed phase droplet. The parameter $C$ is a free parameter that can be defined using curve-fitted monotonic functions on the numerical and experimental data. In [11], it was shown that this single parameter model can account for behavior in a wide range of dimensionless parameter values.
The system of governing equations given by Eqs. (1-2) are then discretized using a Galerkin finite element method using a mixed method [11]. The mixed finite element weak form is given by
@@
where $s$ is the introduced variable for the mixed form that approximates the derivative of the interface distance, that is $\partial h/\partial z$.
The Eqs. (4-6) are solved using continuous Galerkin method with $C^{0}(\Omega)$ elements. The solution algorithm involves moving mesh and calculating the droplet length in a self-consistent way. The algorithm is explained in [22], and has been implemented using the PETSc libraries [23, 24].
The flow becomes highly convective as the droplet interface evolves up to a point when it starts forming the neck. The radial contraction of the interface accelerates faster than the elongation, hence rapidly approaching the singularity. This behavior should be captured accurately to get the precise equilibrium droplet profiles. In addition, the interface is advected with the flow, so that errors in the interface location at one step strongly affect later steps in the flow.
@@ 3 Error estimation
When approaching the singularity, the high-curvature regions quickly develop high stresses. An error in the solution can cause the interface to advect in an incorrect direction, providing an incorrect droplet profile. A coarser mesh must be refined to improve the solution in those regions and get an accurate force balance, leading to more accurate droplet profiles. This refinement can be done adaptively by estimating the true error in the solution using an a posteriori error estimation approach. The foremost goal is to compute the discretization error in the droplet interface evolution.
Since the solution is approximated with $C^{0}(\Omega)$ continuity, the gradients of the solutions are discontinuous across the element boundaries. This discontinuity strongly reflects on the curvature, which can produce erroneous equilibrium droplet profiles. We can use a flux-recovery based approach to drive the adaptive mesh refinement. The flux-recovery approach aims to post-process a smooth gradient from the FE solution and determine the error estimator based on the error in the smoothed and non-smoothed gradients.
In the droplet pinch-off model, we get the velocity ($u$) and the droplet radius ($h$) as the solution. However, the mixed form approach utilizes the approximation of the gradient $\partial h/\partial z$ as a part of the solution. In other words, the mixed variable $s$ already provides a smooth gradient of $h$. If the true gradient is $\partial\bar{h}/\partial z$, the quantity $s$ is assumed to retain better accuracy than $\partial h/\partial z$ since the slope (or flux) $s$ is part of the equilibrium equations and coupled with the curvature derivatives in the momentum equations. The true error norm in the gradient can be written as
@@
The quality of the estimate clearly depends on how good the approximation ‘$s$’ is of the true slope $\partial\bar{h}/\partial z$. Assuming there exists $c\in(0,1)$ such that
@@
Using the triangle inequalities, the bounds on the true error in terms of the error estimate follow directly that
@@
The traditional a posterior error estimation methods like residual-based and flux-based methods use residual in each element, and flux jumps across the elemental interfaces. However, the error bounds achieved in these methods consist of an unknown multiplicative constant.
This approach can also be viewed from the residual-based error estimation perspective. A residual-based a posteriori error estimation approach directly utilizes the finite element solution. Defining the error in the energy norm, one can also derive upper and lower error bounds for the error in the slope ‘$s$’ using the element residuals of the Eq. (6).
Here, the error bounds given in Eq. (9) are used to refine the mesh adaptively. A simple $h$-refinement can be done cyclically:
@@
(1) The elements are marked using the maximum threshold $\bar{\eta}=\lambda\ \max_{K\in\mathcal{T}_{h}}\eta_{K}$, where $\lambda\in[0,1]$. Here, $K$ represents an element from the triangulation $\mathcal{T}_{h}$. This simple strategy allows marking elements $K$ such that $\eta_{K}\geq\bar{\eta}$.
(2) The elements are marked using Dörfler’s method [25], where the elements are marked such that the total error in the marked elements is $\lambda$ times the total error in the entire domain.
@@ 4 Results
The droplet simulations presented in this section are for an 85% glycerol solution with an inlet velocity of 5 mm/s in the background with co-flowing air with 1 m/s velocity. The inlet radius of a nozzle is considered to be 2.5 mm and the radius of the outer capillary tube is 25 mm. The material properties are used from [22].
@@
The error estimation approach is entirely motivated to capture the regions with sharp changes in the mean curvature. Therefore, it is vital to understand the change in the mean curvature to set our expectations for error estimates. Fig. (2) shows a sequence of images showing the change in the mean curvature as the droplet evolves up to the primary pinch-off. Initially, the mean curvature is the largest at the droplet tip. The mean curvature starts increasing in the neck region of the fluid column as the droplet evolves. This is naturally due to the radial shrinkage of the column. Finally, the mean curvature is highest at the pinch-off location when the primary droplet separates from the fluid column. The regions with sharp curvature changes are the top of the neck, the region near the pinch-off, and the droplet tip. Therefore, a higher error measurement is expected in these regions. These images are from the simulation with adaptive mesh refinement for better illustration.
We first report the element-wise errors without performing adaptive mesh refinement. The droplet length and radius are non-dimensionalized using the inlet radius $h_{in}$ in all of the plots. The element-wise error estimate, denoted by $\eta_{K}$, is shown in Fig. (3). The error estimate is plotted for the initial droplet profile and an evolved droplet profile as shown in Fig. (3(a)) and Fig. (3(b)), respectively. The initial droplet profile has the highest curvature gradient closer to the tip. Thus, the highest error is detected at the tip, as the plot reports.
On the other hand, the evolved droplet has a large error at the top of the neck region. This error is associated with the incorrect interface in that region caught in the droplet profile. This inaccuracy leads to a divergence of residuals and simulation failure before reaching the primary pinch-off. The magnified image shows the error magnitude at the tip, which is larger for the last few elements. This suggests that mesh refinement is required to accurately represent a smooth spherical shape of the droplet at the tip.
@@
(a) Non-refined grid.
(b) Adaptively refined grid using Max strategy.
(c) Adaptively refined grid using Dörfler strategy.
@@
Based on the reported errors, the adaptive mesh refinement strategy as explained in the previous section is used for the simulations. As the simulation progresses, we track the error evolution along with the droplet interface. This provides an insight into how the error progresses as the curvature changes. Using the maximum strategy and the Dörfler’s strategy, adaptive mesh refinement is performed when the droplet length is greater than $N$ times the inlet radius, where $N=\{2,3,4,$...\}. The mesh is adaptively refined with $\lambda=0.1$. This means that the elements where the error is larger than 10% of the maximum value are marked and refined. For the Dörfler’s approach, the elements with a cumulative error equal to 90% of the total error are marked and refined. The choice of these criteria is based on the observation in the error evolution for a non-refined grid.
Fig. (4) shows the evolved error superposed on the droplet profile for both no refinement and adaptive refinement. The evolution profiles are shown at the same time stamp. Fig. (4(a)) shows the error plot along the droplet length on a non-refined grid. The top of the neck is a high-velocity region where the upstream flow pushes the droplet axially. However, the radial velocity is much faster than this axial push. Hence as explained before, the discontinuous interface slope (or the gradient) grows quickly. This is visible at the top of the neck where the element-wise error is large. Due to the large error, the droplet interface shows instabilities and hence the simulation diverges. This instability is visible in the droplet profile.
Fig. (4) shows the comparison of the error on non-refined mesh, refined mesh with maximum strategy, and refined mesh with Dörfler’s approach. Compared to the non-refined grid, the error on a refined mesh shows an error reduction in the order of millions. The curvature at the top of the neck region is much smoother, and the interface instabilities are absent. Comparing Fig. (4(b)) and Fig. (4(c)), the Dörfler’s approach requires fewer elements to refine compared to the maximum strategy. The resulting element-wise error in the Dörfler’s strategy case is still on the order of $10^{-6}$.
While comparing only the droplet profile, the non-refined mesh is coarse at the tip, whereas the refined version of the droplet is much smoother. The droplet profiles in mesh with the maximum strategy and the Dörfler’s strategy are almost identical. It is crucial to have an accurate droplet profile to calculate the volume of the pinched-off droplet. The refined version provides accuracy and stability with optimally reduced computational cost.

### v2

@@
Abstract: We present a one-dimensional shear-force-driven droplet formation model with a flux-based error estimator. The model is derived using asymptotic expansion and a front-tracking method to simulate the droplet interface. The model is then discretized using the Galerkin finite element method in the mixed form. However, the solution gradients exhibit large jumps across element boundaries and can grow rapidly due to the highly convective pinch-off process. This leads to an erroneous droplet interface and incorrect curvature. Therefore, the mesh must be sufficiently refined to capture the interface accurately. The mixed form of the governing equation naturally provides smooth interface gradients that can be used to compute the error estimate. The computed error estimate is then used to drive the adaptive mesh refinement algorithm. The efficacy of the error estimator is illustrated by comparing the droplet profiles obtained with adaptive refinement to those obtained with regular refinement. The adaptive mesh refinement approach reduces the computational cost significantly without compromising accuracy.
Keywords: Droplet pinch-off, error estimation, adaptive mesh refinement, front-tracking method, mixed finite element method.
@@ 1 Introduction
The formation of droplets and bubbles is a mature field in physics. These surface-tension-driven phenomena have been deeply explored using interface-tracking and interface-capturing approaches. Historically, the accurate description of free surface flows emerged after Laplace [Laplace1805] and Young [Young1805] explained the role of mean curvature and the surface tension force as a source of instability in 1805. Eggers and Villermaux presented a comprehensive review of the liquid break-up process [EggersVillermaux2008]. Starting with the initial fluid column, the drop becomes heavier by continuously adding fluid. Under an external force like gravity, the instabilities start to grow on the interface. Meanwhile, the surface tension minimizes the surface energy by decreasing the radius of the fluid column. A spherical droplet starts to form at the end of the fluid column hanging from a very thin neck. Eventually, the radius becomes zero and a droplet separates from the initial fluid column. The point of zero radius is called the “pinch-off" point. After the primary pinch-off, the neck recoils due to the unbalanced surface tension. The capillary waves due to this recoil perturb the interface before the tip can collapse back to the top of the fluid cone. This leads to a secondary breakup creating smaller droplets, which are called “satellite drops”.
@@
Inaugural work on the computational modeling of droplet formation was done by Eggers and Dupont [EggersDupont1994]. Their work involved finite difference formulation of the one-dimensional asymptotic model. Ambravaneswaran et al. [AmbravaneswaranWilkesBasaran2002] explored a finite element version of the same model. This model was only explored with the gravitational force as an external force. However, many industrial applications such as atomization and droplet entrainment in annular flow [Inoue_et_al_2021, BernaEscrivaMunozHerranz2015], and microfluidic devices [Cramer2004, Teh2008, Dewandre2020microfluidic] involve droplet formation in a shear-force-driven environment. Using the same asymptotic approach for shear-induced droplets, we considered a one-dimensional model that estimates droplet sizes for an environment with co-flowing fluids [NathawaniKnepleyDropletShear]. For applications with a symmetrical structure of droplet formation, this one-dimensional model with the interface tracking approach is advantageous in terms of computational costs, while still maintaining the accuracy required for comparison with experiments. Nonetheless, this shear-induced droplet model deals with a wide range of external force magnitudes that induce a wide range of interface instability timescales.
@@
The instability leading to the pinch-off point requires a finer mesh in the region of high curvature. As the interface evolves with the progression of numerical simulation, the mesh must be adaptively refined to control discretization errors, especially in the singularity regions. Therefore, a robust error estimation method is essential in designing an adaptive mesh refinement algorithm for our one-dimensional model. The field of a posteriori error estimation has been under continuous and rigorous improvement since the 1980s [GratschBathe2005posteriori]. The preliminary work on error estimation was based on the finite element residuals and the flux jumps across the element boundaries [babuska1978posteriori, babuska1979analysis, babuska1987feedback, kelly1983posteriori]. Since then, many thorough review articles and books have been written on error estimation in finite element analysis; we refer the reader to [ainsworth1997posteriori, becker2001optimal, bangerth2003adaptive]. A thorough review of recent advancements in error estimation is given in [Chamoin_Legoll_Error_Estimation_Review]. In the present work, we show that the mixed finite element formulation of our one-dimensional model naturally provides an effective error estimator for adaptive mesh refinement.
@@
In the following sections, we describe the one-dimensional droplet pinch-off model and explain the need for an error estimator. Then, we explore the flux-based error estimation approach and show that the mixed form naturally provides a smooth flux to compute error bounds. Finally, we illustrate the effectiveness of our error estimator using adaptive mesh refinement.
@@ 2 Problem Description
We consider a fluid column slowly flowing with a constant velocity $u^{d}$ out of a nozzle with radius $h_{in}$. The continuous (or outer) fluid is co-flowing with the dispersed (or inner) fluid column but with a higher velocity $u^{c}$ in a capillary tube with radius $R$. Fig. (1) shows a schematic of this co-flowing fluid scenario. The superscripts $d$ and $c$ represent dispersed phase and continuous phase fluids, respectively.
@@
For the suspended fluid column without a background fluid, we start with the Navier-Stokes equations in cylindrical coordinates. The pinch-off process happens in finite time due to the surface tension forces trying to minimize the surface energy by contracting the interface in the radial direction. The radial contraction is faster than the axial elongation, which allows us to expand the solution variables asymptotically in $r$. Finally, the normal forces are balanced by the surface tension forces, which simplifies the equations to a one-dimensional model.
Similarly, we consider the droplet pinch-off of the fluid column in another co-flowing fluid under the gravitational and shear forces. We consider co-flowing fluids with symmetry preserved in the angular direction. This provides the basis for the one-dimensional approach. The continuous fluid affects the droplet morphology by applying shear forces on the interface between the two fluids. This shear-force effect can be approximated using an asymptotically derived force balance on the interface, yielding the one-dimensional governing equations [NathawaniKnepleyDropletShear]. The full derivation can be found in [NathawaniPhDThesis].
@@
Here, the parameters $\gamma$, $\rho$, $\nu$, $\mu$, and $p$ represent the surface tension coefficient, density, kinematic viscosity, dynamic viscosity, and pressure, respectively. The superscript represents dispersed or continuous phase fluid. $\mathcal{K}$ is the curvature defined by
@@
The quantity (C1)h(C-1)h defines the thickness of the shear layer in the continuous phase flow, which defines how much force is experienced by the dispersed phase droplet. The parameter $C$ is a free parameter that can be defined using curve-fitted monotonic functions on the numerical and experimental data. In [NathawaniKnepleyDropletShear], it was shown that this single parameter model can account for behavior in a wide range of dimensionless parameter values.
The system of governing equations given by Eqs. (1-2) is then discretized using a Galerkin finite element method in the mixed form [NathawaniKnepleyDropletShear]. We will use finite element spaces $\mathcal{U}$ for velocity, $\mathcal{H}$ for radius, and $\mathcal{S}$ for the slope variable, an auxiliary variable introduced in the mixed formulation to approximate the interface slope $\partial h/\partial z$. Now we can express the mixed finite element weak form as
@@
where $u,q\in\mathcal{U}$, $h,v\in\mathcal{H}$, and $s,w\in\mathcal{S}$. Eqs. (4-6) are solved using the continuous Galerkin method with $C^{0}(\Omega)$ elements. The solution algorithm involves a moving mesh and calculates the droplet length in a self-consistent way. The algorithm is explained in [NathawaniKnepleyDropletGravity], and has been implemented using the PETSc libraries [petsc-user-ref, petsc-web-page].
The flow becomes highly convective as the droplet interface evolves up to the point at which the neck begins to form. The radial contraction of the interface accelerates faster than the axial elongation, rapidly approaching the singularity. This behavior must be captured accurately to obtain precise equilibrium droplet profiles. In addition, the interface is advected with the flow, so that errors in the interface location at one step strongly affect later steps in the flow.
@@ 3 Error estimation
When approaching the singularity, the high-curvature regions quickly develop high stresses. An error in the solution can cause the interface to advect in an incorrect direction, yielding an incorrect droplet profile. A coarser mesh must be refined to improve the solution in those regions and obtain an accurate force balance, leading to more accurate droplet profiles. This refinement can be done adaptively by estimating the true error in the solution using an a posteriori error estimation approach. The foremost goal is to compute the discretization error in the droplet interface evolution.
Since the solution is approximated with $C^{0}(\Omega)$ continuity, the solution gradients are discontinuous across the element boundaries. This discontinuity significantly impacts the curvature computation and can produce erroneous equilibrium droplet profiles. We use a flux-recovery-based approach to drive the adaptive mesh refinement. The flux-recovery approach aims to post-process a smooth gradient from the finite element solution and construct the error estimator from the difference between the smoothed and non-smoothed gradients.
In the droplet pinch-off model, the solution consists of the velocity ($u$) and the droplet radius ($h$). The mixed formulation additionally includes an approximation of the gradient $\partial h/\partial z$ as part of the solution, which is denoted by $s$. In other words, the mixed variable $s$ already provides a smooth gradient of $h$. If the true gradient is $\partial\bar{h}/\partial z$, the quantity $s$ is assumed to retain better accuracy than $\partial h/\partial z$ since the slope (or flux) $s$ is part of the equilibrium equations and coupled with the curvature derivatives in the momentum equations. The true error norm in the gradient can be written as
@@
The quality of the estimate clearly depends on how good the approximation $s$ is of the true slope $\partial\bar{h}/\partial z$. Assuming there exists $c\in(0,1)$ such that
@@
Using the triangle inequality, the bounds on the true error in terms of the error estimator follow directly:
@@
The traditional a posteriori error estimation methods, such as residual-based and flux-based methods, use the residual in each element and the flux jumps across the elemental interfaces. However, these error bounds include an unknown multiplicative constant. Our approach can also be viewed from the residual-based error estimation perspective. A residual-based a posteriori error estimation approach directly utilizes the finite element solution. Defining the error in the energy norm, one can also derive upper and lower error bounds for the error in the slope $s$ using the element residuals of Eq. (6).
Here, the error bound given in Eq. (7) are used to refine the mesh adaptively. A simple $h$-refinement can be done cyclically:
@@
(1) The regular refinement, where elements are doubled at each refinement cycle.
(2) The elements are marked using Dörfler’s method [dorfler1996], where the elements are marked such that the total error in the marked elements is $\lambda$ times the total error in the entire domain.
@@ 4 Results
The droplet simulations presented in this section are for an 85% glycerol solution with an inlet velocity of 5 mm/s and a co-flowing air stream at 1 m/s. The nozzle inlet radius is 2.5 mm and the outer capillary tube radius is 25 mm. The material properties are taken from [NathawaniKnepleyDropletGravity].
@@
The error estimation approach is motivated entirely by the need to capture regions with sharp changes in the mean curvature. Therefore, it is important to understand how the mean curvature evolves in order to contextualize the error estimates. Fig. (2) shows a sequence of images illustrating the evolution of mean curvature as the droplet evolves up to the primary pinch-off. The mean curvature is computed using the finite element solution and plotted along the droplet length. The initial mesh is uniform with 50 elements. As the droplet length increases, the mesh is regularly refined by doubling the number of elements as the length increases by a factor of 2 to maintain the accuracy of the solution. The droplet evolves roughly up to 8 times the initial length before the primary pinch-off. Therefore, the mesh is refined up to 800 elements, which 4 levels of refinements. The refinement cycle is defined by the non-dimensionalized droplet length $L/h_{in}$, where $L$ is the droplet length and $h_{in}$ is the inlet radius. We refine the mesh when $L/h_{in}>N$, where $N=\{2,4,6\}$ for the first, second, and third refinement cycle. The last refinement cycle is triggered when $h_{min}<\frac{h_{in}}{50}$, where $h_{min}$ is the minimum droplet radius. This ensures that the mesh is sufficiently refined to capture the pinch-off process. The pinch-off is concluded when $h_{min}<\frac{h_{in}}{100}$.
Initially, the mean curvature is largest at the droplet tip. The mean curvature starts increasing in the neck region of the fluid column as the droplet evolves. This is naturally due to the radial shrinkage of the column. Finally, the mean curvature is highest at the pinch-off location when the primary droplet separates from the fluid column. The regions with sharp curvature changes are the top of the neck, the region near the pinch-off, and the droplet tip. Therefore, a higher error is expected in these regions. One notable observation is that the curvature has a kink at the tip of the droplet. This is due to the choice of initial value of s=10s=-10 at the tip and kept bounded with a large negative value. However, the solution is invariant to the choice of this initial value as long as it is sufficiently negative. Theoretically, the slope at the tip should be negative infinity. The choice of a large negative value is a practical way to approximate this behavior and to keep the simulation stable.
We first report the element-wise errors without performing adaptive mesh refinement. The droplet length and radius are non-dimensionalized using the inlet radius $h_{in}$ in all plots. The element-wise estimated error, denoted by $\eta_{K}$, is shown in Fig. (3). The error is plotted for the initial droplet profile and an evolved droplet profile as shown in Fig. (3(a)) and Fig. (3(b)), respectively. The initial droplet profile has the highest curvature gradient near the tip. Thus, the highest error is detected at the tip, as shown in the figure.
On the other hand, the evolved droplet has a large error at the top of the neck region. This error is associated with an inaccurate interface representation in the neck region. This inaccuracy leads to a divergence of residuals and simulation failure before reaching the primary pinch-off. The magnified image shows the error magnitude at the tip, which is larger for the last few elements. This suggests that mesh refinement is required to accurately capture the smooth spherical shape at the droplet tip.
@@
(a) Regular refinement.
(b) Adaptive refinement using Dörfler strategy.
@@
Based on the reported errors, the adaptive mesh refinement strategy described in the previous section is applied. As the simulation progresses, we track the error evolution along with the droplet interface. This provides insight into how the error evolves as the curvature changes. Using the Dörfler’s strategy, adaptive mesh refinement is performed when the droplet length is greater than $N$ times the inlet radius, where $N=\{2,4,6,$...\}. Similar to the regular refinement strategy, the final refinement is performed when $h_{min}<\frac{h_{in}}{50}$. For the Dörfler’s approach, the elements with a cumulative error equal to 30% of the total error are marked and refined. The choice of these parameters is based on observations from the error evolution on a unrefined grid.
Fig. (4) directly compares regular refinement (left) and adaptive refinement using Dörfler marking (right) at the same evolution stage. The regular strategy, shown in Fig. (4(a)), reduces the error by uniformly adding elements over the full domain. As a result, the error in the neck region is significantly reduced. The largest error is at the pinch-off location, where the curvature changes are highest. Additionally, the error at the top of the neck and at the tip is also apparent. Nonetheless, the error is reduced to the order of $10^{-8}$, which is sufficient to capture the droplet profile accurately. We use this as a reference to compare the Dörfler marking approach, which refines only the elements with the largest error.
Fig. (4(b)) shows the error distribution for the Dörfler marking approach. Similar to the regular refinement, the error is largest at the pinch-off location. The error is reduced to the order of $10^{-2}$, which is smaller than the unrefined case but larger than the regular refinement case. However, the droplet profile is significantly improved compared to the unrefined case, and the instabilities are removed. Moreover, the pinch-off droplet profile is visually indistinguishable from the regular refinement case, indicating that the error level is sufficient to capture the droplet profile accurately.
Comparing Fig. (4(a)) and Fig. (4(b)), both approaches recover a smooth interface and remove the visible instability near the neck. However, a quantitative comparison is also necessary to fully assess the performance of the Dörfler marking approach. Table (1) compares the key quantities of interest and simulation time for a representative regular-refinement run and an AMR run. The quantities are pinch-off location, surface area, volume, and pinch-off time. The AMR case reproduces all quantities with sub $1\%$ inaccuracy compared to the regular-refinement reference, except the pinch-off time which has $1.3174\%$ difference. This is achieved while reducing wall-clock time from $638$ s to $153$ s, which is $76\%$ reduction in computational cost. In other words, this AMR provides a speedup of $4.17\times$, while maintaining close agreement in all quantities of interest. This comparison is evident of the effectiveness of the error estimator in driving the adaptive mesh refinement to capture the droplet profile accurately while significantly reducing computational cost.
The key point is that AMR achieves nearly the same quantities of interest while using far fewer elements. The regular-refinement case reaches 800 elements, whereas the AMR case uses at most 146 elements. This large reduction in mesh size is the main reason for the substantial savings in computational cost.
Table 1: Comparison of regular refinement and AMR for pinch-off metrics and computational cost. Relative differences are computed with respect to the regular-refinement result.
Quantity Regular refinement AMR Relative difference
Non-dimensional pinch-off location 0.7200 0.7186 0.1944%
Pinch-off droplet surface area $7.5548\times 10^{-5}$ $7.5964\times 10^{-5}$ 0.5506%
Pinch-off volume $6.1558\times 10^{-8}$ $6.2043\times 10^{-8}$ 0.7878%
Pinch-off time (s) 0.5769 0.5845 1.3174%
Wall-clock time (s) 638 153 76.02% reduction
Maximum number of elements 800 146 81.75% reduction
The one-dimensional model has been previously used for uncertainty quantification (UQ) analysis of droplet atomization in hybrid rocket combustion [Georgalis2024UQ]. The error estimator can be used to drive the adaptive mesh refinement in the UQ simulations, allowing for more accurate and efficient exploration of the parameter space. This is particularly important in UQ analyses, where a large number of simulations are required to capture the variability in the system. By using the error estimator to refine the mesh adaptively, we can ensure that each simulation is performed with sufficient accuracy while minimizing computational costs.